# Electronic and magnetic properties of graphene-fluorographene nanoribbons: controllable semiconductor-metal transition

### v1

@@ V.2 Spatial structure of $b_{v}$ and $b_{c}$ states
We now seek for a set of orbitals that will allow us to reproduce the $b_{v}$ and $b_{c}$ bands. Our analysis showed that the $b_{v}$ states have a dominant $\mathrm{p_{z}}$ character and minimal hybridization with other orbitals (Fig. 10). Therefore, we can approximate them using a modulated distribution of pure $\mathrm{p_{z}}$ orbitals. In contrast, the $b_{c}$ states are formed through a hybridization of $\mathrm{p_{z}}$ and $\mathcal{A}$ orbitals. As shown in Figs. 9 and 10, the $\mathcal{A}$ states have significant contributions from C-$\mathrm{p_{z}}$, F-$\mathrm{p_{z}}$, C-$\mathrm{s}$, and F-$\mathrm{s}$ orbitals within the fluorographene channels. To understand the nature of these hybridized orbitals, we examine the charge distribution of $\mathcal{A}$ (antibonding) and $\mathcal{B}$ (bonding) states in fluorographene (Fig. 11). Bonding $\mathcal{B}$ states have maximum charge density between atoms, while antibonding $\mathcal{A}$ states exhibit charge density centered on C and F atoms. The antibonding combination, with more nodes, is higher in energy, aligning with the relative energetic positions of $\mathcal{A}$ and $\mathcal{B}$ states in Fig. 9(c). The dominant C-$\mathrm{p_{z}}$ character in $\mathcal{A}$ and F-$\mathrm{p_{z}}$ character in $\mathcal{B}$ are also a consequence of the charge distribution seen in Fig. 11.
Figure 11: (Color online) Schematic illustrating the formation of antibonding $\mathcal{A}$ and bonding $\mathcal{B}$ orbitals in a fluorographene channel. By combining (adding, left) and subtracting (right) atomic $\mathrm{sp}$ orbitals at each $\mathrm{C}$ and $\mathrm{F}$ atoms, the bonding and antibonding $\mathcal{A}$ and $\mathcal{B}$ orbitals are formed. Different colors (grayscale) represent the sign of the phase of the orbital wavefunction.
Figure 12 depicts the minimal orbital basis for a tight binding description of the $b_{v}$ and $b_{c}$ bands of $\alpha\alpha$ and $\alpha\beta$ systems. It comprises the $\mathrm{p_{z}}$ (${\pi}$) orbitals of the graphene channels and the $\mathcal{A}$ orbitals of fluorographene.
Figure 12: (Color online) Orbital basis of the tight-binding Hamiltonian used to simulate the electronic properties of the graphene/fluorographene nanoribbon superlattice. The top and bottom panels show the basis sets for 6-$\alpha\alpha$ and 6-$\alpha\beta$ systems, respectively. Each panel depicts a unit cell containing both graphene and fluorographene regions. The basis set $\mathrm{{\pi}:{\mathcal{A}}}$} comprises ${\pi}$-like orbitals ($\mathrm{p_{z}}$ character) from graphene channels and $\mathrm{{\mathcal{A}}}$ orbitals (primarily $\mathrm{p_{z}}$ character) from fluorographene.
@@ V.3 Effective tight-binding hamiltonian from Wannier90
Our aim is to demonstrate that the $\mathrm{{\pi}}$ and $\mathrm{{\mathcal{A}}}$ orbitals (defined above) are sufficient to reproduce the electronic states around $\mathrm{E_{F}}$ in the $\alpha\alpha$ and $\alpha\beta$ nanoribbons. To achieve this, we employ the Wannier90 package. This software allows us to compute maximally-localized Wannier functions (MLWFs) from a set of Bloch states. This calculation yields a transformation matrix that converts our basis set by combining an algebraic change of basis with an inverse Fourier transform. Furthermore, Wannier90 generates a tight-binding Hamiltonian $\mathcal{H}_{{}_{W}}$ in the MLWF basis using the Kohn-Sham eigenvalues obtained from DFT. With a successful Wannierization, the eigenvalues and eigenvectors of $\mathcal{H}_{{}_{W}}$ exactly reproduce the DFT band structure and Kohn-Sham states within our chosen energy window. This process allows us to obtain an exact tight-binding Hamiltonian for the system.
@@
Figure 13: (Color online) Isosurface plots of the maximally localized Wannier functions (MLWFs) computed using Wannier90. Different colors (grayscale) represent the sign of the wavefunction. The obtained MLWFs closely resemble the initial orbitals proposed for the Wannier calculations (Fig. 12). These results are presented for $6$-$\alpha\alpha$ and 6-$\alpha\beta$ nanoribbons, but similar calculations yield the same MLWF structure for nanoribbons ranging from $4$-$\alpha\beta$ to $8$-$\alpha\beta$, indicating that the MLWFs are independent of system size within this range.
@@
In our case, a successful W90 calculation requires four key inputs: (i) number of MLWFs: to reproduce the $b_{c}$ and $b_{v}$ bands we need $12$ $\mathrm{{\pi}}$ and $12$ $\mathrm{{\mathcal{A}}}$ orbitals for the $6$-$\alpha\alpha$ system, and $13$ $\mathrm{{\pi}}$ and $11$ $\mathrm{{\mathcal{A}}}$ for the $6$-$\alpha\beta$ system (as seen in Fig. 12). Thus, a total of $24$ orbitals for both cases; (ii) energy window for MLWF generation: based on the orbital analysis in Fig. 10, we use [7.5[-7.5[ - 7.5 eV, $13.5$ eV]]]] to encompass all $\mathrm{{\pi}}$ and $\mathrm{{\mathcal{A}}}$ states contributing to the $b_{v}$ and $b_{c}$ bands; (iii) energy window for the tight-binding Hamiltonian: to focus on the $b_{v}$ and $b_{c}$ bands, we select the tighter window [1[-1[ - 1 eV, $2.2$ eV]]]]. This excludes $\mathcal{XY}$ states and higher-energy states with $\mathrm{C-p_{xy}}$ character (Fig. 10); (iv) initial MLWF conditions: We approximate the desired structure using appropriate atomic orbitals as initial seeds for each MLWF. The $\mathrm{{\pi}}$ orbitals are initiated as $\mathrm{C-p_{z}}$ orbitals centered on each carbon atom in the graphene regions while $\mathrm{{\mathcal{A}}}$ as $\mathrm{p_{z}}$ orbitals centered between each C-F pair in fluorographene. The Wannier calculations yielded MLWFs that closely resemble the orbitals proposed in Fig. 11, demonstrating the success of our approach, see (Fig. 13).
@@
Figure 14: (Color online) Comparison of electronic bands obtained from DFT calculations (circles) and the spin-polarized Wannier tight-binding Hamiltonian ($\mathcal{H}_{W}^{\sigma}$, lines). The $\mathcal{H}_{W}^{\sigma}$ eigenvalues accurately reproduce the DFT bands around the Fermi level for all cases. We plot only the bands derived from the ${\pi}$ and $\mathrm{{\mathcal{A}}}$ orbitals (as defined in Fig. 12). For the $n$-$\alpha\beta$ systems, the red and blue lines represent the majority and minority spin bands from $\mathcal{H}_{W}^{\sigma}$, respectively. Since the $\alpha\alpha$ system is non-magnetic, the spin-up and spin-down bands coincide.
@@
The Wannier-based tight-binding Hamiltonian, while exact, contains many parameters. A crucial question for constructing a simplified Anderson-Hubbard model (see next section) is to analyse how to reduce the parameter set while retaining accuracy. Figure 15 addresses this question for the $8$-$\alpha\beta$ nanoribbon. We construct various $\mathcal{H}_{j\sigma}$ Hamiltonians, progressively including interactions up to the fifth nearest neighbor, and compare them to the DFT results. We label these Hamiltonians $\mathcal{H}_{j\sigma}$, where $j$ indicates the farthest neighbor interactions included. It is apparent from the figure that $\mathcal{H}_{5\sigma}$ provides an excellent overall fit to the DFT bands. Importantly, $\mathcal{H}_{j\sigma}$ Hamiltonians with $j<5$ fail to reproduce the semiconductor-to-metal transition observed in $\alpha\beta$ systems (Fig. 6), highlighting the need for up to the fifth-nearest-neighbor interactions to accurately capture this behavior. While $\alpha\alpha$ systems can be reasonably modeled with fewer parameters (second-nearest neighbors), these results underscore the importance of analyzing a larger number of nearest-neighbors to ensure our simplified models to capture the key physics of the $\alpha\beta$ systems.
Figure 15: (Color online) Comparison of accuracy achieved by different simplified tight-binding Hamiltonians ($\mathcal{H}_{j\sigma}$) in describing the DFT results for the $8$-$\alpha\beta$ nanoribbon. $\mathcal{H}_{j\sigma}$ refers to a Hamiltonian where matrix elements beyond the $j$-th nearest neighbor are excluded. Circles represent the DFT data and lines represent the $\mathcal{H}_{j\sigma}$ results. The figure demonstrates that using $\mathcal{H}_{5\sigma}$ is necessary to achieve excellent agreement with the exact DFT results in the $\alpha\beta$ case. Red and blue represent the majority and minority spin bands.
@@ VI Mean field Anderson-Hubbard Model from Wannier calculations
where the operator $c^{{\dagger}}_{j\sigma}$ ($c_{j\sigma}$) creates (annihilates) an electron with spin $\sigma$ at site $j$ (in the corresponding orbital) and $\hat{n}_{j\sigma}=c^{{\dagger}}_{j\sigma}c_{j\sigma}$ is the number operator. Here $\varepsilon_{j}$ is the local energy of site $j$, $t_{jl}$ is the matrix element between $j$ and $l$ sites, and $U_{j}$ is the Hubbard constant at site $j$.
We solve $H_{AH}$ in a mean field description by using the Hartree-Fock method, so the interacting term $\hat{n}_{j\uparrow}\,\hat{n}_{j\downarrow}$ is substituted by $\langle\hat{n}_{j\uparrow}\rangle\,\hat{n}_{j\downarrow}\,+\,\hat{n}_{j\uparrow}\,\langle\hat{n}_{j\downarrow}\rangle$. Thus
@@
where $\varepsilon_{j\sigma}$ is the local energy of spin $\sigma$ that depends of the occupation $\langle\hat{n}_{j\bar{\sigma}}\rangle$ of the opposite spin $\bar{\sigma}$ at the $j$ site, namely
@@
Equation (3) can be used to obtain $U_{j}$, that is
@@
with $\langle\hat{n}_{j\sigma}\rangle=\sum_{\varepsilon_{(\mathrm{k},\sigma)}\leq\mathrm{E_{F}}}|\psi_{(\mathrm{k,\sigma})\,l}|^{2}$. Thus, from $\varepsilon_{j\sigma}$, $\varepsilon_{(\mathrm{k},\sigma)}$, and $\psi_{(\mathrm{k,\sigma})}$, which are the site energy, eigenvalues, and eigenvectors of $\mathcal{H}_{W}^{\sigma}$ one can get an estimate of $U_{j}$ for those sites with non-zero spin polarization. Similarly, the local energy $\varepsilon_{j}$ can also be obtained through Eq. (3), after calculating $U_{j}$,
@@
where $\langle\hat{n}_{j}=\langle\hat{n}_{j\uparrow}\rangle+\langle\hat{n}_{j\downarrow}\rangle\rangle$ is the total electronic occupation at site $j$. A detailed procedure to obtain all the necessary Anderson-Hubbard model parameters is described in the Appendix.
@@
After obtaining the necessary parameters from the analysis of the $6$-$\alpha\alpha$ and $8$-$\alpha\beta$ systems, we now investigate the band structure and magnetic properties of other $n$-$\alpha\alpha$ and $n$-$\alpha\beta$ nanoribbons using the AH model. We first validate our model by contrasting its predictions against DFT calculations for systems up to $10$-$\alpha\alpha$ and $10$-$\alpha\beta$. After this crucial step we extrapolate the AH model and explore larger systems inaccessible to DFT calculations.
@@
Figure 16: (Color online) Validation of the mean field AH model for $n$-$\alpha\alpha$ systems. Upper panels: Comparison of electronic band structures obtained from DFT (open circles) and the AH model (lines) using $\bar{H}_{AH}$ with up to second-nearest neighbor interactions. Results are shown for (a) $6$-$\alpha\alpha$, (b) $8$-$\alpha\alpha$, and (c) $10$-$\alpha\alpha$ systems. Lower panels: Corresponding spin polarization along the graphene channel, calculated using DFT projected density of states (PDOS, red and blue open circles) and $\bar{H}_{AH}$ eigenvectors (black filled circles). Excellent agreement across all systems validates the AH model.
@@
Figure 16 demonstrates the AH model’s accuracy for $6$-$\alpha\alpha$, $8$-$\alpha\alpha$, and $10$-$\alpha\alpha$ nanoribbons. The band structures calculated with the AH model (lines), including up to second-nearest neighbor hopping terms, fairly reproduce the main features of those obtained from DFT (open circles). The AH model reproduces the semiconducting behavior and the band gaps, thus confirming its validity for these systems. While agreement near the Fermi level is excellent, discrepancies arise closer to the Brillouin zone center ($\mathrm{k}=0$). This is to be expected, even for $6$-$\alpha\alpha$, due to our simplified model including only up to second-nearest neighbor interactions. The inclusion of hopping terms up to sixth-nearest neighbors significantly improves the agreement across the entire Brillouin zone and beyond the Fermi level (not shown) [29].
@@
We now explore system sizes inaccessible to DFT. Figure 17 analyzes a $40$-$\alpha\alpha$ nanoribbon—recall from Fig. 5 that the $n$-$\alpha\alpha$ systems exhibit an antiferromagnetic (AF) ground state, with opposite spin alignment across the nanoribbon.
@@
Figure 17: (Color online) Band structure and magnetic properties of the $n$-$\alpha\alpha$ systems. (a) Band structure for a $40$-$\alpha\alpha$ nanoribbon exhibiting a metastable ferromagnetic (FM) state. (b) Band structure for the antiferromagnetic (AF) ground state. (c) Energy difference between AF and FM states, showing the stability of the AF phase and the evolution of its energy gap as a function of $n$.
@@
Figure 17(a) depicts the band structure in the metastable ferromagnetic (FM) state, a revealing metallic behavior. In contrast, the AF ground state (Fig. 17(b)) is semiconducting. Crucially, Fig. 17(c) demonstrates that the AF state is energetically favorable across the studied system sizes (up to $40$-$\alpha\alpha$), although the energy difference ($|\delta E|=|E_{AF}-E_{FM}|$) decreases with increasing size. Figure 17(c) further illustrates how the AF bandgap ($\Delta$) also decreases with increasing $n$.
@@
Figure 18: (Color online) Comparison of DFT spin polarized band structures (red and blue circles) with the mean field AH model predictions (red and blue lines) for $n$-$\alpha\beta$ systems ($n=4,6,8,10$). The AH model accurately captures the key electronic features across all system sizes, validating our approach. Lower panels: corresponding spin polarization along the graphene channel, similar to Fig. 16.
@@
We now consider the $n$-$\alpha\beta$ nanoribbons. As before, we first validate the AH model by comparing it to DFT calculations for $4$-$\alpha\beta$ through $10$-$\alpha\beta$ systems (Fig. 18). We include up to the fifth-nearest neighbors hopping terms based on the $8$-$\alpha\beta$ case (see Appendix). Figure 18 demonstrates an excellent agreement between AH and DFT band structures, specially for large system sizes. Notably, the AH model accurately captures the semiconductor-to-metal transition with increasing system size. The AH model also reproduces the global ferromagnetic state and local spin behavior (lower panels of Fig. 18)—spin polarization oscillates along the channel, as found in DFT calculations. Furthermore, both methods show decreasing spin polarization near the $\beta$ interface with increasing size, converging towards zero magnetic moment at the $\beta$ edge in $10$-$\alpha\beta$. Conversely, the spin polarization at the $\alpha$ edge remains constant, a trend that persists in larger systems (we verified it up to $n=40$).
@@
Figure 19: (Color online) Analysis of magnetic states in larger $\alpha\beta$ systems. (a) Band structure for a $40$-$\alpha\beta$ nanoribbon in the paramagnetic (PM) state. (b) Band structure in the ferromagnetic (FM) state. (c) Energy difference per unit cell ($\delta E=E_{FM}-E_{PM}$) as a function of the number of zigzag chains (n), up to $n=40$. As with smaller systems, spin polarization in (b) is localized around the $\alpha$ interface.

### v2

@@ V.2 Spatial structure of $b_{v}$ and $b_{c}$ states
We now seek a set of orbitals that will allow us to reproduce the $b_{v}$ and $b_{c}$ bands. Our analysis showed that the $b_{v}$ states have a dominant $\mathrm{p_{z}}$ character and minimal hybridization with other orbitals (Fig. 10). Therefore, we can approximate them using a modulated distribution of pure $\mathrm{p_{z}}$ orbitals. In contrast, the $b_{c}$ states are formed through a hybridization of $\mathrm{p_{z}}$ and $\mathcal{A}$ orbitals. As shown in Figs. 9 and 10, the $\mathcal{A}$ states have significant contributions from C-$\mathrm{p_{z}}$, F-$\mathrm{p_{z}}$, C-$\mathrm{s}$, and F-$\mathrm{s}$ orbitals within the fluorographene channels. To understand the nature of these hybridized orbitals, we examine the charge distribution of $\mathcal{A}$ (antibonding) and $\mathcal{B}$ (bonding) states in fluorographene (Fig. 11). Bonding $\mathcal{B}$ states have maximum charge density between atoms, while antibonding $\mathcal{A}$ states exhibit charge density centered on C and F atoms. The antibonding combination, with more nodes, is higher in energy, aligning with the relative energetic positions of $\mathcal{A}$ and $\mathcal{B}$ states in Fig. 9(c). The dominant C-$\mathrm{p_{z}}$ character in $\mathcal{A}$ and F-$\mathrm{p_{z}}$ character in $\mathcal{B}$ are also a consequence of the charge distribution seen in Fig. 11.
Figure 11: Schematic illustrating the formation of antibonding $\mathcal{A}$ and bonding $\mathcal{B}$ orbitals in a fluorographene channel. By combining (adding, left) and subtracting (right) atomic $\mathrm{sp}$ orbitals at each $\mathrm{C}$ and $\mathrm{F}$ atom, the bonding and antibonding $\mathcal{A}$ and $\mathcal{B}$ orbitals are formed. Different colors (grayscale) represent the sign of the phase of the orbital wave function.
Figure 12 depicts the minimal orbital basis for a tight-binding description of the $b_{v}$ and $b_{c}$ bands of $\alpha\alpha$ and $\alpha\beta$ systems. It comprises the $\mathrm{p_{z}}$ (${\pi}$) orbitals of the graphene channels and the $\mathcal{A}$ orbitals of fluorographene.
Figure 12: Orbital basis of the tight-binding Hamiltonian used to simulate the electronic properties of the graphene/fluorographene nanoribbon superlattice. The basis sets for (a) 6-$\alpha\alpha$ and (b) 6-$\alpha\beta$ systems. Each panel depicts a unit cell containing both graphene and fluorographene regions. The basis set $\mathrm{{\pi}:{\mathcal{A}}}$} comprises ${\pi}$-like orbitals ($\mathrm{p_{z}}$ character) from graphene channels and $\mathrm{{\mathcal{A}}}$ orbitals (primarily $\mathrm{p_{z}}$ character) from fluorographene.
@@ V.3 Effective tight-binding Hamiltonian from WANNIER90
Our aim is to demonstrate that the $\mathrm{{\pi}}$ and $\mathrm{{\mathcal{A}}}$ orbitals (defined above) are sufficient to reproduce the electronic states around $\mathrm{E_{F}}$ in the $\alpha\alpha$ and $\alpha\beta$ nanoribbons. To achieve this, we employ the Wannier90 package. This software allows us to compute maximally localized Wannier functions (MLWFs) from a set of Bloch states. This calculation yields a transformation matrix that converts our basis set by combining an algebraic change of basis with an inverse Fourier transform. Furthermore, WANNIER90 generates a tight-binding Hamiltonian $\mathcal{H}_{{}_{W}}$ in the MLWF basis using the Kohn-Sham eigenvalues obtained from DFT. With a successful Wannierization, the eigenvalues and eigenvectors of $\mathcal{H}_{{}_{W}}$ exactly reproduce the DFT band structure and Kohn-Sham states within our chosen energy window. This process allows us to obtain an exact tight-binding Hamiltonian for the system.
@@
Figure 13: Maximally localized Wannier functions (MLWFs) for 6-$\alpha\alpha$ and 6-$\alpha\beta$ nanoribbons, calculated using Wannier90. Isosurfaces are colored (grayscale) according to the sign of the wavefunction. The obtained MLWFs closely resemble the initial guess orbitals (Fig. 12). (a), (c) MLWFs located far from the nanoribbon interfaces; (b), (d) MLWFs at the interfaces. Similar MLWF structures were obtained for nanoribbons ranging from $4$-$\alpha\beta$ to $8$-$\alpha\beta$, indicating that the MLWFs are robust against changes in system size within this range.
@@
In our case, a successful WANNIER90 calculation requires the following four key inputs. (i) Number of MLWFs: To reproduce the $b_{c}$ and $b_{v}$ bands, we need $12$ $\mathrm{{\pi}}$ and $12$ $\mathrm{{\mathcal{A}}}$ orbitals for the $6$-$\alpha\alpha$ system, and $13$ $\mathrm{{\pi}}$ and $11$ $\mathrm{{\mathcal{A}}}$ for the $6$-$\alpha\beta$ system (as seen in Fig. 12). Thus, a total of $24$ orbitals for both cases. (ii) Energy window for MLWF generation: Based on the orbital analysis in Fig. 10, we use [7.5[-7.5[ - 7.5 eV, $13.5$ eV]]]] to encompass all $\mathrm{{\pi}}$ and $\mathrm{{\mathcal{A}}}$ states contributing to the $b_{v}$ and $b_{c}$ bands. (iii) Energy window for the tight-binding Hamiltonian: To focus on the $b_{v}$ and $b_{c}$ bands, we select the tighter window [1[-1[ - 1 eV, $2.2$ eV]]]]. This excludes $\mathcal{XY}$ states and higher-energy states with $\mathrm{C-p_{xy}}$ character (Fig. 10). (iv) Initial MLWF conditions: We approximate the desired structure using appropriate atomic orbitals as initial seeds for each MLWF. The $\mathrm{{\pi}}$ orbitals are initiated as $\mathrm{C-p_{z}}$ orbitals centered on each carbon atom in the graphene, while $\mathrm{{\mathcal{A}}}$ as $\mathrm{p_{z}}$ orbitals centered between each C-F pair in fluorographene. The Wannier calculations yielded MLWFs that closely resemble the orbitals proposed in Fig. 11, demonstrating the success of our approach; see Fig. 13.
@@
Figure 14: Comparison of electronic bands obtained from DFT calculations (circles) and the spin-polarized Wannier tight-binding Hamiltonian ($\mathcal{H}_{W}^{\sigma}$, lines). The $\mathcal{H}_{W}^{\sigma}$ eigenvalues accurately reproduce the DFT bands around the Fermi level for all cases. We plot only the bands derived from the ${\pi}$ and $\mathrm{{\mathcal{A}}}$ orbitals (as defined in Fig. 12). For the $n$-$\alpha\beta$ systems, the red and blue lines represent the majority and minority spin bands from $\mathcal{H}_{W}^{\sigma}$, respectively. Since the $\alpha\alpha$ system is non-magnetic, the spin-up and spin-down bands coincide.
@@
The Wannier-based tight-binding Hamiltonian, while exact, contains many parameters. A crucial question for constructing a simplified Anderson-Hubbard model (see next section) is to analyze how to reduce the parameter set while retaining accuracy. Figure 15 addresses this question for the $8$-$\alpha\beta$ nanoribbon. We construct various $\mathcal{H}_{j\sigma}$ Hamiltonians, progressively including interactions up to the fifth-nearest neighbor, and compare them to the DFT results. We label these Hamiltonians $\mathcal{H}_{j\sigma}$, where $j$ indicates the farthest-neighbor interactions included. It is apparent from the figure that $\mathcal{H}_{5\sigma}$ provides an excellent overall fit to the DFT bands. Importantly, $\mathcal{H}_{j\sigma}$ Hamiltonians with $j<5$ fail to reproduce the semiconductor-to-metal transition observed in $\alpha\beta$ systems (Fig. 6), highlighting the need for up to the fifth-nearest-neighbor interactions to accurately capture this behavior. While $\alpha\alpha$ systems can be reasonably modeled with fewer parameters (second-nearest neighbors), these results underscore the importance of analyzing a larger number of nearest-neighbors to ensure our simplified models capture the key physics of the $\alpha\beta$ systems.
Figure 15: Comparison of accuracy achieved by different simplified tight-binding Hamiltonians ($\mathcal{H}_{j\sigma}$) in describing the DFT results for the $8$-$\alpha\beta$ nanoribbon. $\mathcal{H}_{j\sigma}$ refers to a Hamiltonian where matrix elements beyond the $j$th nearest neighbor are excluded. Circles represent the DFT data and lines represent the $\mathcal{H}_{j\sigma}$ results. The figure demonstrates that using $\mathcal{H}_{5\sigma}$ is necessary to achieve excellent agreement with the exact DFT results in the $\alpha\beta$ case. Red and blue represent the majority and minority spin bands, respectively.
@@ VI Mean field Anderson-Hubbard Model from Wannier calculations
where the operator $c^{{\dagger}}_{j\sigma}$ ($c_{j\sigma}$) creates (annihilates) an electron with spin $\sigma$ at site $j$ (in the corresponding orbital) and $\hat{n}_{j\sigma}=c^{{\dagger}}_{j\sigma}c_{j\sigma}$ is the number operator. Here, $\varepsilon_{j}$ is the local energy of site $j$, $t_{jl}$ is the matrix element between $j$ and $l$ sites, and $U_{j}$ is the Hubbard constant at site $j$.
We solve $H_{AH}$ in a mean-field description by using the Hartree-Fock method, so the interacting term $\hat{n}_{j\uparrow}\,\hat{n}_{j\downarrow}$ is substituted by $\langle\hat{n}_{j\uparrow}\rangle\,\hat{n}_{j\downarrow}\,+\,\hat{n}_{j\uparrow}\,\langle\hat{n}_{j\downarrow}\rangle$. Thus,
@@
where $\varepsilon_{j\sigma}$ is the local energy of spin $\sigma$ that depends of the occupation $\langle\hat{n}_{j\bar{\sigma}}\rangle$ of the opposite spin $\bar{\sigma}$ at the $j$ site, namely,
@@
Equation (3) can be used to obtain $U_{j}$, that is,
@@
with $\langle\hat{n}_{j\sigma}\rangle=\sum_{\varepsilon_{(\mathrm{k},\sigma)}\leq\mathrm{E_{F}}}|\psi_{(\mathrm{k,\sigma})\,l}|^{2}$. Thus, from $\varepsilon_{j\sigma}$, $\varepsilon_{(\mathrm{k},\sigma)}$, and $\psi_{(\mathrm{k,\sigma})}$, which are the site energy, eigenvalues, and eigenvectors of $\mathcal{H}_{W}^{\sigma}$, respectively, one can get an estimate of $U_{j}$ for those sites with nonzero spin polarization. Similarly, the local energy $\varepsilon_{j}$ can also be obtained through Eq. (3), after calculating $U_{j}$,
@@
where $\langle\hat{n}_{j}\rangle=\langle\hat{n}_{j\uparrow}\rangle+\langle\hat{n}_{j\downarrow}\rangle$ is the total electronic occupation at site $j$. A detailed procedure to obtain all the necessary Anderson-Hubbard model parameters is described in the Appendix.
@@
After obtaining the necessary parameters from the analysis of the $6$-$\alpha\alpha$ and $8$-$\alpha\beta$ systems, we now investigate the band structure and magnetic properties of other $n$-$\alpha\alpha$ and $n$-$\alpha\beta$ nanoribbons using the AH model. We first validate our model by contrasting its predictions against DFT calculations for systems up to $10$-$\alpha\alpha$ and $10$-$\alpha\beta$. After this crucial step, we extrapolate the AH model and explore larger systems inaccessible to DFT calculations.
@@
Figure 16: Validation of the mean-field AH model for $n$-$\alpha\alpha$ systems. Upper panels: Comparison of electronic band structures obtained from DFT (open circles) and the AH model (lines) using $\bar{H}_{AH}$ with up to second-nearest-neighbor interactions. Results are shown for (a) $6$-$\alpha\alpha$, (b) $8$-$\alpha\alpha$, and (c) $10$-$\alpha\alpha$ systems. Lower panels: Corresponding spin polarization along the graphene channel, calculated using DFT projected density of states (PDOS, red and blue open circles) and $\bar{H}_{AH}$ eigenvectors (black filled circles). Excellent agreement across all systems validates the AH model.
@@
Figure 16 demonstrates the AH model’s accuracy for $6$-$\alpha\alpha$, $8$-$\alpha\alpha$, and $10$-$\alpha\alpha$ nanoribbons. The band structures calculated with the AH model (lines), including up to second-nearest-neighbor hopping terms, fairly reproduce the main features of those obtained from DFT (open circles). The AH model reproduces the semiconducting behavior and the band gaps, thus confirming its validity for these systems. While agreement near the Fermi level is excellent, discrepancies arise closer to the Brillouin zone center ($\mathrm{k}=0$). This is to be expected, even for $6$-$\alpha\alpha$, due to our simplified model including only up to second-nearest-neighbor interactions. The inclusion of hopping terms up to sixth-nearest neighbors significantly improves the agreement across the entire Brillouin zone and beyond the Fermi level (not shown) [29].
@@
We now explore system sizes that are inaccessible to DFT. Figure 17 analyzes a $40$-$\alpha\alpha$ nanoribbon—recall from Fig. 5 that the $n$-$\alpha\alpha$ systems exhibit an antiferromagnetic (AF) ground state, with opposite spin alignment across the nanoribbon.
@@
Figure 17: Band structure and magnetic properties of the $n$-$\alpha\alpha$ systems, calculated within the mean-field AH model. (a) Band structure for a $40$-$\alpha\alpha$ nanoribbon exhibiting a metastable ferromagnetic (FM) state. (b) Band structure for the antiferromagnetic (AF) ground state. (c) Energy difference between AF and FM states, showing the stability of the AF phase and the evolution of its energy gap as a function of $n$. DFT results are shown for comparison (squares).
@@
Figure 17(a) depicts the band structure in the metastable ferromagnetic (FM) state, a revealing metallic behavior. In contrast, the AF ground state [Fig. 17(b)] is semiconducting. Crucially, Fig. 17(c) demonstrates that the AF state is energetically favorable across the studied system sizes (up to $40$-$\alpha\alpha$), although the energy difference ($|\delta E|=|E_{AF}-E_{FM}|$) decreases with increasing size. Figure 17(c) further illustrates how the AF bandgap ($\Delta$) also decreases with increasing $n$.
@@
Figure 18: Comparison of DFT spin-polarized band structures (red and blue circles) with the mean-field AH model predictions (red and blue lines) for $n$-$\alpha\beta$ systems ($n=4,6,8,10$). The AH model accurately captures the key electronic features across all system sizes, validating our approach. Lower panels: corresponding spin polarization along the graphene channel, similar to Fig. 16.
@@
We now consider the $n$-$\alpha\beta$ nanoribbons. As before, we first validate the AH model by comparing it to DFT calculations for $4$-$\alpha\beta$ through $10$-$\alpha\beta$ systems (Fig. 18). We include up to the fifth-nearest-neighbor hopping terms based on the $8$-$\alpha\beta$ case (see the Appendix). Figure 18 demonstrates an excellent agreement between AH and DFT band structures, specially for large system sizes. Notably, the AH model accurately captures the semiconductor-to-metal transition with increasing system size. The AH model also reproduces the global ferromagnetic state and local spin behavior (lower panels of Fig. 18)—spin polarization oscillates along the channel, as found in DFT calculations. Furthermore, both methods show decreasing spin polarization near the $\beta$ interface with increasing size, converging towards zero magnetic moment at the $\beta$ edge in $10$-$\alpha\beta$. Conversely, the spin polarization at the $\alpha$ edge remains constant, a trend that persists in larger systems (we verified it up to $n=40$).
@@
Figure 19: Analysis of magnetic states in larger $\alpha\beta$ systems. (a) Band structure for a $40$-$\alpha\beta$ nanoribbon in the paramagnetic (PM) state. (b) Band structure in the ferromagnetic (FM) state. (c) Energy difference per unit cell ($\delta E=E_{FM}-E_{PM}$) as a function of the number of zigzag chains ($n$), up to $n=40$. DFT results are shown for comparison (squares). As with smaller systems, spin polarization in (b) is localized around the $\alpha$ interface.